A hailstone with diameter of 5.6 fell. It weighed about 0.018 lb/in compared to the normal 0.033 lb/in for ice. About how heavy was this hailstone?
Answer by stanbon(75887)

A hailstone with diameter of 5.6 in fell. It weighed about 0.018 lb/in^3 compared to the normal 0.033 lb/in^3 for ice. About how heavy was this hailstone?
---------------
Volume of the hailstone:
V = (4/3)pi*r^3 = (4/3)pi*2.8^3 = 91.95 in^3
------
Weight:: 91.95*0.018 = 1.66 lbs
-----------------------------------------
Cheers,
Stan H.
===================